# Forecast constraints on the axion-photon coupling from interstellar medium heating

### v1

@@ IV Heating rate induced by forced resonance
Since the ratio $r_{\nu}=\nu/\omega_{p}$ of interstellar media given in Table 1 is in the region $10^{-13}\lesssim r_{\nu}\lesssim 10^{-5}$, we have chosen three different values of $\nu$ in this range in Fig. 1. In the left panel, which corresponds to $\dot{C}=10^{-27}$ erg cm-3 s-1 and $B_{0}=10^{-5}$ G, we obtain $g_{\rm max}=1.9\times 10^{-16}$ GeV-1 for the resonance mass $m_{\phi}=\omega_{p}$ with $r_{\nu}=10^{-13}$ [case (a)]. For $r_{\nu}=10^{-9}$ [case (b)] and $r_{\nu}=10^{-5}$ [case (c)], we have $g_{\rm max}=1.9\times 10^{-14}$ GeV-1 and $g_{\rm max}=1.9\times 10^{-12}$ GeV-1, respectively. As we observe in the left panel of Fig. 1, the upper limits of $g_{\rm max}$ for these three cases are tighter than the observational bound derived from polarization measurements of the MWD stars. Away from the resonance axion mass, however, the values of $g_{\rm max}$ are much larger than those obtained at $m_{\phi}=\omega_{p}$. For the choices of $\dot{C}$ and $B_{0}$ given in the left panel of Fig. 1, the upper limits of $g_{\rm max}$ derived in the mass ranges $m_{\phi}\ll\omega_{p}$ and $m_{\phi}\gg\omega_{p}$ are much weaker than those computed at $m_{\phi}=\omega_{p}$.
@@
If we were to obtain tighter observational bounds on $\dot{C}$, Eq. (59) shows that $g_{\rm max}$ gets smaller for any mass $m_{\phi}$. This property also holds for $B_{0}$, where the observational finding of interstellar media with $B_{0}$ larger than $10^{-5}$ G results in smaller values of $g_{\rm max}$. In the right panel of Fig. 1, we show $g_{\rm max}$ versus $m_{\phi}$ for $\dot{C}=10^{-31}$ erg cm-3 s-1 and $B_{0}=1$ G. As we see in Eq. (59), the coupling $g_{\rm max}$ for given $m_{\phi}$ and $\nu$ is $10^{-7}$ times as small as the one in the left panel. As a result, even the mass range outside the resonance point can enter the parameter space constrained by the MWD measurements. This tendency is more significant for increasing values of $\nu$, even if $g_{\rm max}$ at $m_{\phi}=\omega_{p}$ gets larger. In case (c), for example, the mass range satisfying the bound $g\leq 5.4\times 10^{-12}$ GeV-1 corresponds to $3.5\times 10^{-14}~{}{\rm eV}\leq m_{\phi}\leq 2.8\times 10^{-9}~{}{\rm eV}$. Thus, for smaller $\dot{C}$ and larger $B_{0}$, it is possible to obtain tighter constraints on $g_{\rm max}$ with wider parameter spaces of $m_{\phi}$.
@@
While the data of $\dot{C}$ shown in Table 1 correspond to those of the typical interstellar media, there may be possibilities to find some data of particular regions in the Milky Way galaxy or dwarf galaxies which result in tighter limits on $\dot{C}$ with larger $B_{0}$. We also note that the plasma frequency $\omega_{p}$ can vary for a wider range than those given in Table 1 ($10^{-12}~{}{\rm eV}\lesssim m_{\phi}\lesssim 10^{-10}~{}{\rm eV}$). Then, we may be able to obtain more stringent constraints on $g_{\rm max}$ with the wider region of $m_{\phi}$ in future observations.
@@ V Conclusions
In this paper, we proposed a new method of probing the axion-photon coupling through the heating of plasmas in interstellar media. We assumed the presence of an interstellar magnetic field aligned in a particular direction, whose typical value is in the range $B_{0}=10^{-5}~{}{\rm G}\sim 10^{-3}$ G. We also considered a nonrelativistic thermalized plasma of electrons and heavy ions. The observed cooling rate of interstellar media $\dot{C}$ as well as the electron number density $n_{e}$ and the thermal temperature $T$ are given in Table 1. The latter two data can translate to the plasma frequency $\omega_{p}$ and the friction constant $\nu$, all of which satisfy the condition $\omega_{p}\gg\nu$ with $10^{-12}~{}{\rm eV}\lesssim\omega_{p}\lesssim 10^{-10}~{}{\rm eV}$. Although the precise values of $B_{0}$ have been unknown for the data shown in Table 1, we performed forecast constraints on the axion-photon coupling constant $g$ in preparation for upcoming observations of the interstellar plasma.
@@
In Sec. IV, we showed that the axion-photon coupling leads to the heating of interstellar media through the resonant energy conversion from axions to photons. In the direction parallel to the external magnetic field, the heating rate $\dot{Q}$ balances the energy damping rate $\dot{Q}_{\nu}$, which represents the energy release from photons to the plasma induced by the friction $\nu$. Since $\dot{Q}$ should not exceed the observed cooling rates $\dot{C}$ of interstellar media, it is possible to put upper bounds on the axion-photon coupling $g$. At the resonance point, $g$ is constrained as Eq. (58). For the typical interstellar magnetic field strength $B_{0}=10^{-5}$ G, we found that $g_{\rm max}$ at the resonance point ($m_{\phi}=\omega_{p}$) can be more stringent than the limit constrained by the MWD measurements. As we see in the left panel of Fig. 1, which corresponds to $B_{0}=10^{-5}$ G and $\dot{C}=10^{-27}$ erg cm-3 s-1, $g_{\rm max}$ is much larger than the MWD bound for $m_{\phi}$ away from $\omega_{p}$. However, for larger $B_{0}$ and smaller $\dot{C}$, even the mass range outside the resonance point can enter the region constrained by the MWD measurements. This behavior is clearly seen in the right panel of Fig. 1.
We have thus shown that observations of the interstellar magnetic field along with the cooling rates of plasmas will offer an interesting possibility for constraining the axion-photon coupling $g$. Since the resonance is most efficient for $m_{\phi}$ around $\omega_{p}$ that ranges between $10^{-12}$ eV and $10^{-10}$ eV, we may be able to obtain new bounds on $g$ for the mass region like $10^{-15}~{}{\rm eV}\lesssim m_{\phi}\lesssim 10^{-7}~{}{\rm eV}$ with upcoming data. This will bring a new perspective for scrutinizing the properties of light axions as a candidate for dark matter.
@@ Acknowledgements
We thank Shoichi Yamada for useful discussions. This work was supported by the Grant-in-Aid for Scientific Research Fund of the JSPS Nos. 20H05854 (TF), 22K03642 (ST), 23K03424 (TF), and Waseda University Special Research Project No. 2024C-474 (ST).
@@
$E_{0}\simeq 2\sqrt{n_{e}T}\simeq 5.1\times 10^{-8}\,{\rm eV}^{2}\left(\frac{n_{e}}{0.1\,{\rm cm}^{-3}}\right)^{1/2}\left(\frac{T}{10^{4}\,{\rm K}}\right)^{1/2}\,.$ (65)
The background magnetic field of order $10^{-5}$ G corresponds to $B_{0}=1.95\times 10^{-7}$ eV2. For $n_{e}\approx 0.1\,{\rm cm}^{-3}$ and $T\approx 10^{4}$ K, $E_{0}$ is of a similar order to $B_{0}\approx 10^{-5}$ G.
@@
$g\ll 2.1\times 10^{9}\,{\rm GeV}^{-1}\left(\frac{m_{\phi}}{10^{-11}\,{\rm eV}}\right)\left(\frac{B_{0}}{10^{-5}\,{\rm G}}\right)^{-1}\left(\frac{\rho_{\rm DM}}{0.3~{}{\rm GeV\,cm}^{-3}}\right)^{1/2}\left(\frac{n_{e}}{0.1~{}{\rm cm}^{-3}}\right)^{-1/2}\left(\frac{T}{10^{4}~{}{\rm K}}\right)^{-1/2}\,.$ (66)
@@
In comparison to Eq. (26), the presence of the magnetic field ${\bm{B}}_{0}$ gives rise to terms proportional to $\omega_{c}$ associated with the quasi-circular motion in the $(y,z)$ plane. Combing Eq. (67) with Eq. (68), we obtain the following fourth-order differential equations

### v2

@@ IV Heating rate induced by forced resonance
Since the ratio $r_{\nu}=\nu/\omega_{p}$ of interstellar media given in Table 1 is in the region $10^{-13}\lesssim r_{\nu}\lesssim 10^{-7}$, we have chosen three different values of $\nu$ in this range in Fig. 1. In the left panel, which corresponds to $\dot{C}=10^{-27}$ erg cm-3 s-1 and $B_{0}=10^{-6}$ G, we obtain $g_{\rm max}=6.0\times 10^{-16}$ GeV-1 for the resonance mass $m_{\phi}=\omega_{p}$ with $r_{\nu}=10^{-13}$ [case (a)]. For $r_{\nu}=10^{-10}$ [case (b)] and $r_{\nu}=10^{-7}$ [case (c)], we have $g_{\rm max}=1.9\times 10^{-14}$ GeV-1 and $g_{\rm max}=6.0\times 10^{-13}$ GeV-1, respectively. As we observe in the left panel of Fig. 1, the upper limits of $g_{\rm max}$ for these three cases are tighter than the observational bound derived from polarization measurements of the MWD stars. Away from the resonance axion mass, however, the values of $g_{\rm max}$ are much larger than those obtained at $m_{\phi}=\omega_{p}$. For the choices of $\dot{C}$ and $B_{0}$ given in the left panel of Fig. 1, the upper limits of $g_{\rm max}$ derived in the mass ranges $m_{\phi}\ll\omega_{p}$ and $m_{\phi}\gg\omega_{p}$ are much weaker than those computed at $m_{\phi}=\omega_{p}$. This reflects the fact that the resonance width is given by $\Delta m_{\phi}=\nu\ll\omega_{p}$, in which case the resonance occurs in a narrow region around $m_{\phi}=\omega_{p}$.
@@
If we were to obtain tighter observational bounds on $\dot{C}$, Eq. (59) shows that $g_{\rm max}$ gets smaller for any mass $m_{\phi}$. This property also holds for $B_{0}$, where the observational finding of interstellar media with $B_{0}$ larger than $10^{-6}$ G results in smaller values of $g_{\rm max}$. In the right panel of Fig. 1, we show $g_{\rm max}$ versus $m_{\phi}$ for $\dot{C}=10^{-31}$ erg cm-3 s-1 and $B_{0}=1$ G. As we see in Eq. (59), the coupling $g_{\rm max}$ for given $m_{\phi}$ and $\nu$ is $10^{-8}$ times as small as the one in the left panel. As a result, even the mass range outside the resonance point can enter the parameter space constrained by the MWD measurements. This tendency is more significant for increasing values of $\nu$, even if $g_{\rm max}$ at $m_{\phi}=\omega_{p}$ gets larger. In case (c), for example, the mass range satisfying the bound $g\leq 5.4\times 10^{-12}$ GeV-1 corresponds to $1.1\times 10^{-14}~{}{\rm eV}\leq m_{\phi}\leq 9.0\times 10^{-11}~{}{\rm eV}$. Thus, for smaller $\dot{C}$ and larger $B_{0}$, it is possible to obtain tighter constraints on $g_{\rm max}$ with wider parameter spaces of $m_{\phi}$ away from the resonance point even if the resonance band is narrow.
@@
While the data of $\dot{C}$ shown in Table 1 correspond to those of the typical interstellar media, there may be possibilities to find some data of particular regions in the Milky Way galaxy or dwarf galaxies which result in tighter limits on $\dot{C}$ with larger $B_{0}$. Upcoming observations such as Vera Rubin Observatory and future Subaru HSC (Hyper Suprime-Cam) observations are expected to find other gas-rich faint dwarf galaxies like Leo T, which may obtain the cooling rate smaller than the current limit. We also note that the plasma frequency $\omega_{p}$ can vary for a wider range than $10^{-12}~{}{\rm eV}$. Then, we may be able to obtain more stringent constraints on $g_{\rm max}$ with the wider region of $m_{\phi}$ in future observations.
@@ V Conclusions
In this paper, we proposed a new method of probing the axion-photon coupling through the heating of plasmas in interstellar media. We assumed the presence of an interstellar magnetic field aligned in a particular direction, whose typical value is of order $B_{0}=10^{-6}$ G. We also considered a nonrelativistic thermalized plasma of electrons and heavy ions. The observed cooling rate of interstellar media $\dot{C}$ and the thermal temperature $T$ are given in Table 1. For the typical electron number density $n_{e}\simeq 10^{-3}~{}{\rm cm}^{-3}$, the plasma frequency $\omega_{p}$ and the friction constant $\nu$ can be computed by Eq. (3), so that $\nu\ll\omega_{p}\simeq 10^{-12}~{}{\rm eV}$ for all of these interstellar media. Although the precise values of $B_{0}$ have been unknown for the data shown in Table 1, we performed forecast constraints on the axion-photon coupling constant $g$ in preparation for upcoming observations of the interstellar plasma.
@@
In Sec. IV, we showed that the axion-photon coupling leads to the heating of interstellar media through the resonant energy conversion from axions to photons. In the direction parallel to the external magnetic field, the heating rate $\dot{Q}$ balances the energy damping rate $\dot{Q}_{\nu}$, which represents the energy release from photons to the plasma induced by the friction $\nu$. Since $\dot{Q}$ should not exceed the observed cooling rates $\dot{C}$ of interstellar media, it is possible to put upper bounds on the axion-photon coupling $g$. At the resonance point, $g$ is constrained as Eq. (58). For the typical interstellar magnetic field strength $B_{0}=10^{-6}$ G, we found that $g_{\rm max}$ at the resonance point ($m_{\phi}=\omega_{p}$) can be more stringent than the limit constrained by the MWD measurements. As we see in the left panel of Fig. 1, which corresponds to $B_{0}=10^{-6}$ G and $\dot{C}=10^{-27}$ erg cm-3 s-1, $g_{\rm max}$ is much larger than the MWD bound for $m_{\phi}$ away from $\omega_{p}$. However, for larger $B_{0}$ and smaller $\dot{C}$, even the mass range outside the resonance point can enter the region constrained by the MWD measurements. This behavior is clearly seen in the right panel of Fig. 1.
We have thus shown that observations of the interstellar magnetic field along with the cooling rates of plasmas will offer an interesting possibility for constraining the axion-photon coupling $g$. Since the resonance is most efficient for the mass $m_{\phi}$ around $\omega_{p}$, we may be able to obtain new bounds on $g$ for the mass region like $10^{-15}~{}{\rm eV}\lesssim m_{\phi}\lesssim 10^{-9}~{}{\rm eV}$ with upcoming data. This will bring a new perspective for scrutinizing the properties of light axions as a candidate for dark matter.
@@ Acknowledgements
We are grateful Shoichi Yamada for fruitful discussions. We also thank Zihui Wang for useful correspondence. This work was supported by the Grant-in-Aid for Scientific Research Fund of the JSPS Nos. 20H05854 (TF), 22K03642 (ST), 23K03424 (TF), and Waseda University Special Research Project No. 2024C-474 (ST).
@@
$E_{0}\simeq 2\sqrt{n_{e}T}\simeq 5.1\times 10^{-9}\,{\rm eV}^{2}\left(\frac{n_{e}}{10^{-3}\,{\rm cm}^{-3}}\right)^{1/2}\left(\frac{T}{10^{4}\,{\rm K}}\right)^{1/2}\,.$ (65)
The background magnetic field of order $10^{-6}$ G corresponds to $B_{0}=1.95\times 10^{-8}$ eV2. For $n_{e}\approx 10^{-3}\,{\rm cm}^{-3}$ and $T\approx 10^{4}$ K, $E_{0}$ is of a similar order to $B_{0}\approx 10^{-6}$ G.
@@
$g\ll 2.1\times 10^{10}\,{\rm GeV}^{-1}\left(\frac{m_{\phi}}{10^{-12}\,{\rm eV}}\right)\left(\frac{B_{0}}{10^{-6}\,{\rm G}}\right)^{-1}\left(\frac{\rho_{\rm DM}}{0.3~{}{\rm GeV\,cm}^{-3}}\right)^{1/2}\left(\frac{n_{e}}{10^{-3}~{}{\rm cm}^{-3}}\right)^{-1/2}\left(\frac{T}{10^{4}~{}{\rm K}}\right)^{-1/2}\,.$ (66)
@@
In comparison to Eq. (26), the presence of the magnetic field ${\bm{B}}_{0}$ gives rise to terms proportional to $\omega_{c}$ associated with the quasi-circular motion in the $(y,z)$ plane. Combining Eq. (67) with Eq. (68), we obtain the following fourth-order differential equations